A football coach starts at the 10-yard line and walks to the 40-yard line. He turns around and walks to the 0 yard line. He turns once more and walks back to the 50 yard line. He turns around once more and walks back to the 25 yard line. Thus completing his 10 minutes of pacing. The coach's overall distance is ____.

A. 15yd, east B.15 yd, west C. 15yd
D.50yd, east E. 50yd, west F. 50yd
G. 60yd, east H. 60yd west I. 60yd
J. 85yd east K. 85yd west L. 85yd
M. 145 yd east N. 145yd west O. 145yd

or P. none of these
1
Expert's answer
2012-10-01T10:41:44-0400
1. A football coach starts at the 10-yard line and walks to the 40-yard line.

So, the first part of his displacement is

D1 = 40 - 10 = 30 yards.

2. He turns around and walks to the 0 yard line.

Coach starts from the 40-yard line and walks to the 0 yard line, so his displacement is

D2 = 40 - 0 = 40 yards.

3. He turns once more and walks back to the 50 yard line.

Coach starts from the 0-yard line and walks to the 50-yard line, so his displacement is

D3 = 50 - 0 = 50 yards.

4. He turns around once more and walks back to the 25 yard line.

Coach starts from the 50-yard line and walks to the 25-yard line, so his displacement is

D4 = 50 - 25 = 25 yards.

The total distance walked is

D = D1 + D2 + D3 + D4 = 30 + 40 + 50 + 25 = 145 yards.

So, correct answer is O.